OpenStudy (anonymous):

5x – y = 16 5x + y = 14 I need to solve for x and y

OpenStudy (anonymous):

i assume i tansfer it to be y=-5+16 and y=5-14?

OpenStudy (anonymous):

and then solve to the intercept

OpenStudy (lgbasallote):

you actually use systems of linear equations...either by substitution or by eliminations..but is the instruction to use slope-intercept?

OpenStudy (anonymous):

it just says to solve by graphing

OpenStudy (anonymous):

Solving Systems of Linear Equations by Graphing and the Substitution Method

OpenStudy (lgbasallote):

substitution method...you need to substitute one equation into the other..i dont know what graphing method is but i surely know sub :DDD

OpenStudy (anonymous):

I think I graph it till both lines meet and thats my point and answer for the x and y